# Artinian groups of large cardinality

### v1

@@ Abstract.
A group is Artinian if there is no infinite strictly descending chain of subgroups. Ol’shanskii has asked whether there are Artinian groups of arbitrarily large cardinality. We reduce this problem to an analogous question, regarding universal algebras, asked by Jónsson in the 1960s. We provide Artinian groups of cardinality $\aleph_{n}$ for each natural number $n$. We also give a consistent strong negative answer to the question of Ol’shanskii (from a large cardinal assumption) as well as a consistent positive answer. Thus, answers to the questions of Ol’shanskii and Jónsson are independent of set theory.
@@ 1. Introduction
For a fixed infinite cardinal $\lambda$ it is easy to see that exactly one of two possibilities can arise. It could be the case that there exist Artinian algebras of all cardinalities having signature of size at most $\lambda$. In the other case there exists a cardinal $\kappa$ such that algebras of size $\kappa$ or larger and signature of size at most $\lambda$ are never Artinian, but there are Artinian algebras having signature of size at most $\lambda$ at all cardinals below $\kappa$. In the latter case we write $\textbf{artin}(\lambda)=\kappa$ and in the former $\textbf{artin}(\lambda)=\infty$. Theorem A (1) tells us that there are Artinian groups at all cardinals greater than 00 and below $\textbf{artin}(\aleph_{0})$, and no Artinian algebras of countable signature at or above $\textbf{artin}(\aleph_{0})$. So, determining the value of $\textbf{artin}(\aleph_{0})$ simultaneously answers the question of Ol’shanskii and that of Jónsson.
@@
For each infinite cardinal $\lambda$ the equality $\textbf{fr}(\lambda)=\textbf{artin}(\lambda)$ holds.
@@ 2. Proof of Theorem A
We argue why Theorem A (2) holds. Whaley has shown that if there is a finite signature Artinian algebra of cardinality $\aleph_{\alpha}$ then there also exists one of the next higher cardinality $\aleph_{\alpha+1}$ [15, Theorem 2.3], and a Artinian algebras of finite signature exist at $\aleph_{n}$ for each $n<\omega$. Thus we apply the equivalence of (a) and (b).
@@ 2. $(*)_{5}$
    There is a simple torsion-free group of cardinality $\kappa$ having no infinite free subset.
@@
Suppose for contradiction that $\{g_{n}\}_{n<\omega}$ is an infinite free subset of $G$. For each $k<\omega$ we let $H_{k}=\langle\{g_{n}\}_{n\geq k}\rangle$. So $\{H_{k}\}_{k<\omega}$ is a strictly descending chain of subgroups of $G$, and none of the $H_{k}$ can be cyclic (since the infinite set $\{g_{n}\}_{n\geq k}$ is free in $H_{k}$ and a cyclic group clearly cannot have an infinite free subset). Using Lemma 2.5, inductively conjugate this sequence as before to a strictly descending sequence $K_{0}>K_{1}>\cdots$ with $K_{k}=\langle C_{k}\rangle$ with $C_{k}=f(X_{k})$ for some $X_{k}\subseteq\Omega$, and $X_{k}\not\subseteq G_{\alpha}$ for all $\alpha$. Obtain the same contradiction. Thus $G$ has the desired properties and (1) holds.

### v2

@@ Abstract.
A group is Artinian if there is no infinite strictly descending chain of subgroups. Ol’shanskii has asked whether there are Artinian groups of arbitrarily large cardinality. We show that this problem is essentially the same as an analogous question, regarding universal algebras, asked by Jónsson in the 1960s. We further show that these problems are the same as the so-called free subset problem. As a result, one can have a consistent strong negative answer (from a large cardinal assumption) as well as a consistent positive answer.
@@ 1. Introduction
For a fixed infinite cardinal $\lambda$ it is easy to see that exactly one of two possibilities can arise. It could be the case that there exist Artinian algebras of all cardinalities having signature of size at most $\lambda$. In the other case there exists a cardinal $\kappa$ such that algebras of size $\kappa$ or larger and signature of size at most $\lambda$ are never Artinian, but there are Artinian algebras having signature of size at most $\lambda$ at all cardinals below $\kappa$. In the latter case we write $\textbf{artin}(\lambda)=\kappa$ and in the former $\textbf{artin}(\lambda)=\infty$. Theorem A (1) tells us that there are Artinian groups at all cardinals greater than 00 and below $\textbf{artin}(\aleph_{0})$, and no Artinian algebras of countable signature at or above $\textbf{artin}(\aleph_{0})$. So, determining the value of $\textbf{artin}(\aleph_{0})$ simultaneously answers the question of Ol’shanskii and that of Jónsson. To establish further results we move into the setting of the free subset problem.
@@
For each infinite cardinal $\lambda$ one has $\textbf{fr}(\lambda)=\textbf{artin}(\lambda)$.
@@ 2. Proof of Theorem A
We argue why Theorem A (2) holds. Whaley has shown that if there is a finite signature Artinian algebra of cardinality $\aleph_{\alpha}$ then there also exists one of the next higher cardinality $\aleph_{\alpha+1}$ [15, Theorem 2.3], and Artinian algebras of finite signature exist at $\aleph_{n}$ for each $n<\omega$. Thus we apply the equivalence of (a) and (b).
@@ 2. $(*)_{5}$
    There is a simple torsion-free group of cardinality $\kappa$ having no infinite free subset, and in which every strictly descending chain of noncyclic subgroups is finite.
@@
Suppose for contradiction that $\{g_{n}\}_{n<\omega}$ is an infinite free subset of $G$. For each $k<\omega$ we let $H_{k}=\langle\{g_{n}\}_{n\geq k}\rangle$. So $\{H_{k}\}_{k<\omega}$ is a strictly descending chain of subgroups of $G$, and none of the $H_{k}$ can be cyclic (since the infinite set $\{g_{n}\}_{n\geq k}$ is free in $H_{k}$ and a cyclic group clearly cannot have an infinite free subset). Using Lemma 2.5, inductively conjugate this sequence as before to a strictly descending sequence $K_{0}>K_{1}>\cdots$ with $K_{k}=\langle C_{k}\rangle$ with $C_{k}=f(X_{k})$ for some $X_{k}\subseteq\Omega$, and $X_{k}\not\subseteq G_{\alpha}$ for all $\alpha$. Obtain the same contradiction. From our proof, every strictly descending chain of noncyclic subgroups in $G$ is finite. Thus $G$ has the desired properties and (1) holds.